OpenStudy (anonymous):

Owen scored eighty-five points on a one hundred-point test and twenty-six on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.

OpenStudy (anonymous):

85/100 = 17/20 or 26/30 = 13/15 Which one has the larger fraction of the total points ?

OpenStudy (anonymous):

13/15

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

tyvm :)

OpenStudy (anonymous):

so your answer is 26/30

OpenStudy (strawberry17):

So you have 85/100 for the first test and 26/30 for the second test. Reduce fractions to find a common denominator easier: 85 5 17 --- ÷ -- = -- 100 5 20 26 2 13 -- ÷ -- = -- 30 2 15 You would want to get a common denominator of 60. So numerator and denominator of 17/20 by 3 and multiply the numerator and denominator of 13/15 by 4. 85/100 = 17/20 = 26/30 = 13/15 =

OpenStudy (anonymous):

[3.04]Owen scored eighty-five points on a one hundred-point test and twenty-six on a thirty-point quiz. On which did he get a larger fraction of the total points correct? Show your work.